OpenStudy (anonymous):

Marcia a lab technician needs three liters of an 8% saline solution. She has a 5% and 9% solution in the lab stock room. How many liters of the 5% and 9% solution should she mix. (slove using linear systems)

OpenStudy (anonymous):

i hope she should mix 2.25lt of 9% solution and .75lt of 5% solution

OpenStudy (anonymous):

What equations did you use to get that

OpenStudy (anonymous):

@harsh314 ?? I know that's the right answer but how??

OpenStudy (anonymous):

ok you need the following equations salinity of the required solution is 8% which means 8lt in100lt now the volume of solution is 3lt so the amount of salinity in 3lt is 24/100(i) let us take x lt from 5% saline solution and y lt from other one now salinity in x=5x/100 salinity in y=9y/100 total salinity =(5x+9y)/100(ii) from (i) and (ii) 5x+9y=24 and as you know the net amount is 3lt so x+y=3 solve them you get the answers...........

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (anonymous):

my pleasure ................